OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 A. (3, 1) B. (-1, 3) C. (-1, -3) D. (1, 3)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Can you explain why it is D?

OpenStudy (anonymous):

-2x + y = 1 y=1+2x put this y in another equation -4x + 1+2x = -1 -2x=-2 x=1 put x=1 any one of equation u will get y =3

OpenStudy (anonymous):

Thanks man.